one person with a skiploader requires 12 h to remove a large quantity of earth. a second, larger skiploader can remove the same amount of earth in 4 h. how long would it take to remove the earth with both skiploaders working together?
Answer by htmentor(1343)

The first person's removal rate is 1 per 12 hours (e.g. 1 ton per 12 hrs)
The second person's removal rate is 1 per 4 hours
Working together their rate in equation form is:
1/12 + 1/4 = 1/12 + 3/12 = 4/12 = 1/3
This says that the removal rate working together is 1 (ton) per 3 hrs
Taking the reciprocal we get 3 hrs per 1 ton
Ans: 3 hours